If h (θ) = cos2θ, find h(0), h (1/4π), h(1/2π)


1
Expert's answer
2022-06-06T16:02:52-0400
"h(0)=\cos^2(0)=1^2=1"

"h(\pi/4)=\cos^2(\pi/4)=(\dfrac{\sqrt{2}}{2})^2=\dfrac{1}{2}"

"h(\pi/2)=\cos^2(\pi/2)=(0)^2=0"